In preparing an infusion of the drug trastuzumab, a pharmacist adds 20 ml of bacteriostatic water for injection to a vial containing 440 mg of the mAb. The resulting solution contains trastuzumab, 21 mg/ml. Using the information: Trastuzumab loading dose: 4 mg/kg with 90 min infusion

(a) Calculate the milliliters of solution required for the loading dose for a 165 lb. Patient.

(b) The pharmacist then transfers the calculated volume into an infusion bag containing 250 ml of sodium chloride injection. Calculate the quantity of trastuzumab administered in mg/min.

(c) Calculate the drip rate in drops/min, using an administered set that delivers 20 drops/ml

ONLY ANSWER B.. and C if you want. I have A-- 14.3
Answer #1

Given, In preparing an infusion of the drug trastuzumab, a pharmacist adds 20 ml of bacteriostatic water for injection to a vial containing 440 mg of the mAb.
The resulting solution contains , 21 mg/ml.

Trastuzumab loading dose: 4 mg/kg with 90 min infusion

(a) patient weight= 165 lb

Now, 1 lb = 0.454 kg
165 lb= 0.454×165 kg
=74.91 kg
Loading dose for the patient= dose per kg × patient weight
=4 mg/kg × 74.91 kg
=299.64 mg

Now, resulting solution contain trastuzumab 21 mg/ml

So, 21 mg trastuzumab drug present in 1 ml
1 mg trastuzumab drug present in=1/21 ml
299.64 mg trastuzumab drug present in=(1/21)×299.64 ml
=14.3 ml
Therefore, 14.3 ml of solution required for the loading dose for 165 lb patient.
(b) given, pharmacist transfer the calculated volume that is 14.3 ml into an infusion bag containing 250 ml of sodium chloride injection.
So, total volume in bag = 250+14.3 =264.3 ml
Now, 14.3 ml solution contain 299.64 mg trastuzumab.
This 299.64 mg trastuzumab now present in 264.3 ml solution as 14.3 ml mixed with 250 ml solution.
Now volume of solution in the infusion bag = 264.3 ml
Infusion time = 90 min
So, infusion rate = total volume in ml in the bag/infusion time in min
=264.3 ml/90 min
=2.94 ml/min
Now, 264.3 ml solution contain 299.64 mg trastuzumab
1 ml solution contain =299.64/264.3 mg trastuzumab
2.94 ml solution contain=(299.64/264.3)×2.94 mg trastuzumab
=3.33 mg trastuzumab
So, infusion rate = 2.94 ml/min
=3.33mg/min

Hence, 2.94 ml solution contain 3.33 mg trastuzumab drag

Therefore, the quantity of trastuzumab administered = 3.33 mg/min.